determine the equation of the line through (7,3), which is parallel to the line defined by 2x+3y=4
Answer by mananth(16946)

2 x 3 y = 4
Find the slope of this line
make y the subject
3 y = -2 x + 4
Divide by 3
y = - 2/3 x + 4/3
Compare this equation with y=mx+b
slope m = - 2/3
The slope of a line parallel to the above line will be the same
The slope of the required line will be - 2/3
m= - 2/3 ,point ( 7 , 3 )
Find b by plugging the values of m & the point in
y=mx+b
3 = -14/ 3 + b
b= 23/3
m= - 2/3
Plug value of the slope and b in y = mx +b
The required equation is y = - 2/3 x + 23/ 3

m.ananth@hotmail.ca